OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 13 2x + y = 2

OpenStudy (angela210793):

x - y = 13---> x=13+y 2x + y = 2 replace in here 2(13+y)+y=2 26+2y+y=2 3y=-24 y=-8 x=13-8=5

OpenStudy (anonymous):

How about showing some work your self????

OpenStudy (anonymous):

\[-2x + 2y = -26 \] \[2x + 2y = 2\] 4y = -24 y = -6